# Scaling limits for fractional polyharmonic Gaussian fields

### v3

@@ 4.2. Scaling limit in Besov, Sobolev and Hölder spaces
This is essentially [11, Theorem 2.30]. There a local version of the theorem is given. The global version presented here is obtained by choosing $U=\mathbb{R}^{d}$, $\hat{\Omega}\subset K_{1}\subset K_{2}\subset\ldots$ such that already $K_{1}$ is far larger than $\hat{\Omega}$, $k_{1}=k_{2}=\ldots=0$ and observing that, for functions with uniformly compact support, the local and global Besov spaces agree. Finally, the assertion that $\int_{\mathbb{R}^{d}}x^{\underline{m}}g_{j}(x)\,\mathrm{d}x=0$ is stated in [11, Equation (2.2)]. ∎
@@
Since $\operatorname{supp}(\Theta_{h}*\tilde{f}(x-\cdot))\subset\tilde{\tilde{\Omega}}$, we use the Poincaré inequality (Lemma A.2) to deduce that the $\|\cdot\|_{\dot{H}^{-s}_{h}(\tilde{\tilde{\Omega}}_{h})}$-norm is bounded by a multiple of the (inomogenous) $\|\cdot\|_{H^{-s}_{h}(h\mathbb{Z}^{d})}$-norm. Indeed, if we fix another bounded domain $\Omega^{*}\Supset\tilde{\tilde{\Omega}}$ and a cut-off function $\eta\in C_{c}^{\infty}(\Omega^{*})$ that is equal to 1 in $\tilde{\tilde{\Omega}}$ and let $\eta_{h}$ be its restriction to $h\mathbb{Z}^{d}$, then for any $g_{h}\colon h\mathbb{Z}^{d}\to\mathbb{R}$ that is supported in $\tilde{\tilde{\Omega}}_{h}$ we have
@@
$\leq C\left(\int_{\left(-\frac{\pi}{h},\frac{\pi}{h}\right)^{d}}(1+M_{h}(\xi)^{2})^{s}|\mathcal{F}_{h}[\eta_{h}](\xi)|^{2}\,\mathrm{d}\xi\right)\left(\int_{\left(-\frac{\pi}{h},\frac{\pi}{h}\right)^{d}}|\mathcal{F}_{h}[j_{h}](\xi)|\,\mathrm{d}\xi\right)^{2}$
$\qquad+C\left(\int_{\left(-\frac{\pi}{h},\frac{\pi}{h}\right)^{d}}|\mathcal{F}_{h}[\eta_{h}](\xi)|^{2}\,\mathrm{d}\xi\right)\left(\int_{\left(-\frac{\pi}{h},\frac{\pi}{h}\right)^{d}}(1+M_{h}(\xi)^{2})^{s}|\mathcal{F}_{h}[j_{h}](\xi)|\,\mathrm{d}\xi\right)^{2}$
which, by the Poincaré inequality on $\Omega_{h}^{*}$ and the fact that $\mathcal{F}_{h}[\eta_{h}]$ decays faster than any polynomial (uniformly in $h$), is bounded by $C\|j_{h}\|_{\dot{H}^{s}_{h}(\Omega^{*}_{h})}$.
@@
Step 1. Consequences of Theorem 1.2. Let $k\in\mathbb{N}$, with $k>s+\frac{d}{2}$, be arbitrary and take $\Theta$ to be a product of one-dimensional B-splines of order $k$ as in [18, Section 1.9.4], i.e.,
@@
(4.18) $\max_{\begin{subarray}{c}y,\,y^{\prime}\in h\mathbb{Z}^{d}\\ |y-y^{\prime}|y_{\infty}\leq kh\end{subarray}}|f_{h}(y)-f_{h}(y^{\prime})|\leq Ch^{\alpha}\|I_{h}f_{h}\|_{C^{0,\,\alpha}(\mathbb{R}^{d})}.$
@@
Let us take $x\in{h_{*}}\mathbb{Z}^{d}$. Then $I_{h_{*}}f_{h_{*}}\restriction_{x+(0,{h_{*}}/2)^{2}}$ is a polynomial of degree at most $k-1$ in each variable, which depends precisely on the values of $f_{h_{*}}$ in $x+\left(-\frac{{h_{*}}k}{2},-\frac{{h_{*}}(k+1)}{2}\right)\cap h\mathbb{Z}^{d}$. The space of polynomials of degree at most $k-1$ in each variable is an $\mathbb{R}$-vector space of dimension exactly $k^{d}$. The same holds true for the space of functions from $x+\left(-\frac{{h_{*}}k}{2},-\frac{{h_{*}}(k+1)}{2}\right)\cap{h_{*}}\mathbb{Z}^{d}$ to $\mathbb{R}$. This means that $I_{h_{*}}f_{h_{*}}$ induces a linear map between two finite-dimensional vector spaces of the same dimension. This map has trivial kernel (as follows for example from [24, Corollary 4.62] together with an induction on $d$) and so it is in fact an isomorphism.
@@
then $\lim_{M\to\infty}\lim_{h\to 0}\mathbb{P}(\mathcal{E}_{M})=1$. On the other hand, (4.18) implies that, on the event $\mathcal{E}_{M}$, we have
@@
$\lim_{M\to\infty}\lim_{h\to 0}\mathbb{P}\left(\max_{y\in\mathbb{R}^{d}}I_{h}\varphi_{h}(y)\leq\max_{x\in\Omega_{h}}\varphi_{h}(x)\leq\max_{y\in\mathbb{R}^{d}}I_{h}\varphi_{h}(y)+CMh^{\alpha}\right)=1,$
@@ Appendix B Fractional Gaussian Fields via eigenfunctions
For $s^{\prime}\in\{-s,0,s\}$ the estimate (B.1) follows directly from the definition. We next claim that (B.1) holds whenever $-s\leq s^{\prime}\leq s$. To see this, we adapt the argument in [21, Corollary 1]. Namely, we take first $0<s^{\prime}<s$, consider $(-\Delta)^{s}$ restricted to functions in $\dot{H}^{s}(\Omega)$, and let $((-\Delta)^{s})_{N}^{s^{\prime}/s}$ be its (spectral) $\frac{s^{\prime}}{s}$-th power. Explicitly,
@@
Claim (i). By the Hilbert-space-valued version of Kolmogorov’s two series theorem (see e.g. [14, Corollary on p. 386]), the series

### v4

@@ 4.2. Scaling limit in Besov, Sobolev and Hölder spaces
This is essentially [12, Theorem 2.30]. There a local version of the theorem is given. The global version presented here is obtained by choosing $U=\mathbb{R}^{d}$, $\hat{\Omega}\subset K_{1}\subset K_{2}\subset\ldots$ such that already $K_{1}$ is far larger than $\hat{\Omega}$, $k_{1}=k_{2}=\ldots=0$ and observing that, for functions with uniformly compact support, the local and global Besov spaces agree. Finally, the assertion that $\int_{\mathbb{R}^{d}}x^{\underline{m}}g_{j}(x)\,\mathrm{d}x=0$ is stated in [12, Equation (2.2)]. ∎
@@
Since $\operatorname{supp}(\Theta_{h}*\tilde{f}(x-\cdot))\subset\tilde{\tilde{\Omega}}$, we use the Poincaré inequality (Lemma A.2) to deduce that the $\|\cdot\|_{\dot{H}^{-s}_{h}(\tilde{\tilde{\Omega}}_{h})}$-norm is bounded by a multiple of the (inhomogenous) $\|\cdot\|_{H^{-s}_{h}(h\mathbb{Z}^{d})}$-norm. Indeed, if we fix another bounded domain $\Omega^{*}\Supset\tilde{\tilde{\Omega}}$ and a cut-off function $\eta\in C_{c}^{\infty}(\Omega^{*})$ that is equal to 1 in $\tilde{\tilde{\Omega}}$ and let $\eta_{h}$ be its restriction to $h\mathbb{Z}^{d}$, then for any $g_{h}\colon h\mathbb{Z}^{d}\to\mathbb{R}$ that is supported in $\tilde{\tilde{\Omega}}_{h}$ we have
@@
$\leq C\left(\int_{\left(-\frac{\pi}{h},\frac{\pi}{h}\right)^{d}}(1+M_{h}(\xi)^{2})^{s/2}|\mathcal{F}_{h}[\eta_{h}](\xi)|^{2}\,\mathrm{d}\xi\right)^{2}\left(\int_{\left(-\frac{\pi}{h},\frac{\pi}{h}\right)^{d}}|\mathcal{F}_{h}[j_{h}](\xi)|^{2}\,\mathrm{d}\xi\right)$
$\qquad+C\left(\int_{\left(-\frac{\pi}{h},\frac{\pi}{h}\right)^{d}}|\mathcal{F}_{h}[\eta_{h}](\xi)|\,\mathrm{d}\xi\right)^{2}\left(\int_{\left(-\frac{\pi}{h},\frac{\pi}{h}\right)^{d}}(1+M_{h}(\xi)^{2})^{s}|\mathcal{F}_{h}[j_{h}](\xi)|^{2}\,\mathrm{d}\xi\right)$
which, by the Poincaré inequality on $\Omega_{h}^{*}$ and the fact that $\mathcal{F}_{h}[\eta_{h}]$ decays faster than any polynomial (uniformly in $h$), is bounded by $C\|j_{h}\|_{\dot{H}^{s}_{h}(\Omega^{*}_{h})}^{2}$.
@@
Step 1. Consequences of Theorem 1.2. Let $k\in\mathbb{N}$, with $k>s+\frac{d}{2}$, be arbitrary and take $\Theta$ to be a product of one-dimensional B-splines of order $k$ as in [19, Section 1.9.4], i.e.,
@@
(4.18) $\max_{\begin{subarray}{c}y,\,y^{\prime}\in h\mathbb{Z}^{d}\\ |y-y^{\prime}|_{\infty}\leq kh\end{subarray}}|f_{h}(y)-f_{h}(y^{\prime})|\leq Ch^{\alpha}\|I_{h}f_{h}\|_{C^{0,\,\alpha}(\mathbb{R}^{d})}.$
@@
Let us take $x\in{h_{*}}\mathbb{Z}^{d}$. Then $I_{h_{*}}f_{h_{*}}\restriction_{x+(0,{h_{*}}/2)^{d}}$ is a polynomial of degree at most $k-1$ in each variable, which depends precisely on the values of $f_{h_{*}}$ in $x+\left(-\frac{{h_{*}}k}{2},\frac{{h_{*}}(k+1)}{2}\right)^{d}\cap h\mathbb{Z}^{d}$. The space of polynomials of degree at most $k-1$ in each variable is an $\mathbb{R}$-vector space of dimension exactly $k^{d}$. The same holds true for the space of functions from $x+\left(-\frac{{h_{*}}k}{2},\frac{{h_{*}}(k+1)}{2}\right)^{d}\cap{h_{*}}\mathbb{Z}^{d}$ to $\mathbb{R}$. This means that $I_{h_{*}}f_{h_{*}}$ induces a linear map between two finite-dimensional vector spaces of the same dimension. This map has trivial kernel (as follows for example from [24, Corollary 4.62] together with an induction on $d$) and so it is in fact an isomorphism.
@@
then $\lim_{M\to\infty}\lim_{h\to 0}\mathbb{P}_{h}(\mathcal{E}_{M})=1$. On the other hand, (4.18) implies that, on the event $\mathcal{E}_{M}$, we have
@@
$\lim_{M\to\infty}\lim_{h\to 0}\mathbb{P}_{h}\left(\max_{y\in\mathbb{R}^{d}}I_{h}\varphi_{h}(y)\leq\max_{x\in\Omega_{h}}\varphi_{h}(x)\leq\max_{y\in\mathbb{R}^{d}}I_{h}\varphi_{h}(y)+CMh^{\alpha}\right)=1,$
@@ Appendix B Fractional Gaussian Fields via eigenfunctions
For $s^{\prime}\in\{-s,0,s\}$ the estimate (B.1) follows directly from the definition. We next claim that (B.1) holds whenever $-s\leq s^{\prime}\leq s$. To see this, we adapt the argument in [22, Corollary 1]. Namely, we take first $0<s^{\prime}<s$, consider $(-\Delta)^{s}$ restricted to functions in $\dot{H}^{s}(\Omega)$, and let $((-\Delta)^{s})_{N}^{s^{\prime}/s}$ be its (spectral) $\frac{s^{\prime}}{s}$-th power. Explicitly,
@@
Claim (i). By the Hilbert-space-valued version of Kolmogorov’s two series theorem (see e.g. [15, Corollary on p. 386]), the series